# Cosmic (super)strings with a time-varying tension

### v1

@@ 3.3 Loop decays
In order to maintain the scaling regime, the network must lose energy at the same rate as the background. This is done by emission of loops. These loops then decay into axions, massive matter or gravitational waves, depending on the origin of the defect. We mentioned the discrepancies in the literature regarding these decay channels in the introduction, and here we will focus in local, Nambu-Goto strings. Simulations Blanco-Pillado:2013qja suggest that around $90\%$ of the energy of the network is released into small ultra-relativistic loops that are red-shifted away. The other fraction $\mathcal{F}_{\alpha}\simeq 0.1$ of the energy is deposited into small loops of length that is a fraction $\alpha$ of the correlation length:
$\ell(t_{i})=\alpha L(t_{i})=\alpha\xi t_{i}\,,$ (34)
where in the second equality we have assumed the scaling regime of the VOS. Local cosmic string loops decay through gravitational radiation. [^12] Global strings can also decay by radiating massless Nabu-Goldstone bosons, with $\frac{dE_{\text{\tiny{GW}}}}{dt}\propto v^{2}$, where $v$ is the vev of the corresponding scalar field. In the frequency regime of interest, the power of emission is independent of their length, and reads
$\frac{dE_{\text{\tiny{GW}}}}{dt}=\Gamma G\mu^{2}\,,$ (35)
@@
$\ell(t)=\alpha t_{i}-\Gamma G\mu(t-t_{i})\,.$ (36)
Let us assume that the network loses energy at a rate Kibble:1984hp
$\frac{d\rho}{dt}=\tilde{c}\bar{v}\frac{\rho}{L}\,,$ (37)
@@
$\mathcal{F}_{\alpha}\frac{d\rho}{dt}=\mu\gamma\ell\frac{dn}{dt}\to\frac{dn}{dt}=\frac{\mathcal{F}_{\alpha}}{\gamma}\frac{\tilde{c}\bar{v}}{\alpha}\frac{1}{L^{4}}\,.$ (38)
where $\gamma$ is a Lorentz factor which takes into account that $\ell$ is the proper length of the loop, while the energy is computed in the rest frame of the mother string. The loop production rate during scaling therefore features a behaviour $dn/dt=A/t^{4}$ during scaling, where $A$ depends on the background Blanco-Pillado:2017oxo ; Blanco-Pillado:2013qja and can be inferred from simulations.
@@ 4 The effect of a time-dependent tension
$\mu=m_{s}^{2}=\left(\frac{M_{P}}{4\pi}\right)^{2}\frac{\sqrt{g_{s}}}{\mathcal{V}},$ (39)
and the extra-dimensional volume $\mathcal{V}$ is measured in stringy units of $(2\pi\sqrt{\alpha^{\prime}})^{6}$. In phenomenologically interesting scenarios, such as volume kination or tracker epochs Apers:2024ffe , the volume can vary with time, and thus induce a time-dependent string tension. In both cases the tension evolves as $\mu\sim t^{-m}$, with $m=1$ and $m=2/3$ for kination and the tracker respectively. The goal of this section will be to extend the original VOS model with a time-dependent string tension, and also discuss the evolution of loops131313 An early attempt to study the VOS with varying tension Yamaguchi:2005gp imposed the string tension to depend on coordinate, rather than worldsheet time, in the Nambu-Goto action. This breaks worldsheet reparametrisations and leads to the wrong equations of motion (see also Emond:2021vts ; Conlon:2024uob ). It would be interesting to adapt the methods of Martins:2018dqg to the equations presented in this article.. We shall see that allowing for a varying tension entails a rich phenomenology, making scaling far from obvious. The goal of this article is to argue that a scaling solution is nevertheless achieved, and we will make some comments on the timescales for its stability to the percolation phenomenon recently discussed in Conlon:2024uob . Towards the end of the section we will also extend the analysis to the population of loops, whose distribution is crucial for the computation of gravitational wave spectra.
@@ 4.1 Equations of motion and analytical aspects
$S=-\int d^{2}\zeta\,\mu\left(X^{\nu}(\zeta)\right)\sqrt{-\gamma}.$ (40)
@@
$\frac{1}{\mu\left(X^{\nu}\right)\sqrt{-\gamma}}\partial_{a}\left(\mu\left(X^{\nu}\right)\sqrt{-\gamma}\gamma^{ab}g_{\rho\sigma}X^{\sigma}_{\,,b}\right)-\frac{1}{2}\partial_{\rho}g_{\lambda\sigma}\gamma^{ab}X^{\lambda}_{\,,a}X^{\sigma}_{\,,b}-\frac{\partial_{\rho}\mu\left(X^{\nu}\right)}{\mu\left(X^{\nu}\right)}=0,$ (41)
where the dependence on $X^{\nu}(\zeta)$ is left implicit. Eq. (41) was previously derived in Emond:2021vts ; Conlon:2024uob , and differs from the equation in Yamaguchi:2005gp only by the last term, which arises precisely through the dependence of the tension on the embedding coordinates. From now onwards, we also assume that the tension only depends on time, that is $\mu\equiv\mu\left(X^{0}(\zeta)\right)$. As in the case of constant tension, we choose a gauge where $\dot{X}\cdot X^{\prime}=0$ (and $\gamma_{01}=\gamma_{10}=0$), and identify conformal and worldsheet time, $X^{0}(\zeta)=\tau$. Notice, however, that this is only possible after the equations of motion have been written down out, as doing so in the action would have lead to a wrong result for (41). Defining the energy density of a string as in (23), the equation of motion (41) for $\rho=0$ obtains an additional term with respect to the constant tension case
$\dot{\epsilon}+\left(2\frac{\dot{a}}{a}+\frac{\dot{\mu}}{\mu}\right)\varepsilon\dot{x}^{2}=0.$ (42)
@@
$\ddot{x}_{i}+\left(2\frac{\dot{a}}{a}+\frac{\dot{\mu}}{\mu}\right)(1-\dot{x}^{2})\dot{x}_{i}-\frac{1}{\varepsilon}\left(\frac{x_{i}^{{}^{\prime}}}{\varepsilon}\right)^{{}^{\prime}}=0,$ (43)
@@
$2\mathcal{H}^{\prime}=2\mathcal{H}+\frac{\dot{\mu}}{\mu}.$ (44)
@@
The interplay between varying tension and Hubble expansion leads to a rich phenomenology. For instance, as explained in Emond:2021vts , for the special case $\mu\sim a(t)^{-2}$ the action becomes scale-invariant, and the equations of motion reproduce those of a constant tension string in flat space-time. Even more striking is the case when the RHS of (44) is negative, since the loops grow faster than the scale factor Conlon:2024uob ; the strings behave as if they were placed in a contracting universe. It is this case that we intend to study in this article [^15] Note that the universe is still undergoing expansion, so the dynamics of the network will be different to that of cosmic strings in an actual contracting universe, studied in Avelino:2002hx ; Avelino:2002xy ..
@@ 4.1.1 The circular loop
To be more concrete about our claims, let us find a solution for the microscopic equations of motion in an arbitrary power-law cosmology $a\sim t^{2/n}$ with $\mu\sim t^{-m}$. Note that this includes the constant tension case $m=0$, but we are not aware of this solution having been reported elsewhere in the literature (although its qualitative features are understood). It is well known that, when $\textbf{x}^{\prime}\cdot\dot{\textbf{x}}=0$, the equations of motion in Eqs. (42) and (43) are not independent. To solve the condition $\textbf{x}^{\prime}\cdot\dot{\textbf{x}}=0$ we focus on the circular loop:
$\textbf{x}=R(\tau)L(\cos(\zeta^{1}/L),\sin(\zeta^{1}/L),0)\,,$ (45)
@@
$R_{max}(\tau)\sim\left(\frac{\tau}{\tau_{0}}\right)^{-q/2}\,,$ (47)
@@ 4.1.2 The momentum parameter: analytical modelling
The momentum parameter $k(v)$ plays a crucial role in the averaged equations of motion, and in this section we would like to gain some understanding on its behaviour from a microscopic point of view. As will be clear in the next section, a crucial aspect of achieving a scaling regime (when $m\,n>4$) is that the momentum parameter switches sign when crossing the threshold $v^{2}=\frac{1}{2}$. Let us provide some evidence in support of this claim. To do so, we study the evolution of small scale structure with varying tension. Let us follow Vilenkin:2000jqa and study linear perturbations with respect to the static, straight string:
$\textbf{x}=\textbf{c}\zeta^{1}\,,$ (48)
@@
$\delta\ddot{x}_{i}+\left(2\frac{\dot{a}}{a}+\frac{\dot{\mu}}{\mu}\right)\delta\dot{x}_{i}-\delta x_{i}^{\prime\prime}=0\,.$ (49)
@@
$\delta x=\textbf{A}_{k}\left(\frac{\tau_{0}}{\tau}\right)^{\nu}J_{\nu}(k\tau)e^{ik\zeta^{1}}+\textbf{B}_{k}\left(\frac{\tau_{0}}{\tau}\right)^{\nu}Y_{\nu}(k\tau)e^{ik\zeta^{1}}\,,$ (50)
@@
$\lim_{k\tau\to\infty}J_{\nu}(k\tau)=\sqrt{\frac{2}{\pi k\tau}}\cos\left(z-\frac{\nu\pi}{2}-\frac{\pi}{4}\right)\,,$ (51)
@@
$\langle(1-\dot{\textbf{x}}^{2})\dot{\textbf{x}}\cdot\textbf{u}\rangle=\langle\frac{R}{a}\frac{1}{\epsilon^{2}}\textbf{x}^{\prime\prime}\cdot\dot{\textbf{x}}\rangle\,.$ (52)
@@
$\langle\frac{R}{a}\frac{1}{\epsilon^{2}}\textbf{x}^{\prime\prime}\cdot\dot{\textbf{x}}\rangle=\frac{1}{16a}\textbf{A}_{k}^{2}k^{2}\left(\left(\frac{\tau_{0}}{\tau}\right)^{2\nu}J_{\nu}(k\tau)^{2}-\frac{1}{2^{\nu}\Gamma(1+\nu)^{2}}\right)\,,$ (53)
where $\Gamma(x)$ is Euler’s gamma function and we have averaged, for simplicity, from an initial time $\tau=0$. The crucial point here is that for subhorizon perturbations, using Eq. (51), the first term dominates for $02\nu+1<$ 0, which is equivalent to $mn<4$. Otherwise the second term dominates and the contribution to the momentum parameter is negative - we take this as suggestive evidence of a flip of sign in the momentum parameter for the VOS model when $mn>4$ and therefore for the existence of scaling solutions in this regime, as explained below. We believe this is so even if our solutions do not fully capture the complicated microscopic configuration of the long strings in a scaling solution. In particular, the microscopic equations cannot take into account the loss of energy into loops by interactions of the strings, which is a key ingredient for scaling to occur.
@@ 4.2 Long strings
In this section, we examine how the modified equations of motion (42)-(43) will impact the evolution and properties of the network, focusing on the population of the long strings.
@@ 4.2.1 Network evolution
From (42), the energy density $\rho\sim\frac{E}{a^{3}}$ of strings then obeys the equation
$\frac{{\rm d}\rho}{{\rm d}t}=-2H(1+v^{2})\rho+\frac{1}{\mu}\frac{{\rm d}\mu}{{\rm d}t}(1-v^{2})\rho$ (54)
where we have switched to the FRW time coordinate $t$. Notice that, in the non-relativistic limit $(v\rightarrow 0)$, the energy density redshifts as $\rho\sim\mu(t)/a(t)^{2}$, while in the ultra-relativistic limit $(v\rightarrow 1)$ it behaves as radiation, $\rho\sim 1/a(t)^{4}$. From the definition of the characteristic length $\rho=\frac{\mu}{L^{2}}$, one can turn (54) into an equation for $L$. For long strings, one also has to take into account energy loss into loops, which can be modelled as
$\frac{{\rm d}\rho}{{\rm d}t}=-\tilde{c}v\frac{\rho}{L}.$ (55)
Putting everything together, (54)-(55) can be combined to give
$2\frac{{\rm d}L}{{\rm d}t}=2HL(1+v^{2})+\tilde{c}v+\frac{L}{\mu}\frac{{\rm d}\mu}{{\rm d}t}v^{2}.$ (56)
Following the same steps, one can use (43) to derive a new equation for $v$, namely
$\frac{{\rm d}v}{{\rm d}t}=(1-v^{2})\left[\frac{k(v)}{L}-\left(2H+\frac{1}{\mu}\frac{{\rm d}\mu}{{\rm d}t}\right)v\right].$ (57)
The difference between these and those of Martins:2018dqg were explained in footnote 13. Here, one can already see a crucial, qualitative difference with respect to the constant tension case: if $02H+\frac{1}{\mu}\frac{{\rm d}\mu}{{\rm d}t}<$ 0, the RHS of (57) cannot be negative for any $v$ satisfying $k(v)>0$, forbidding the existence of a fixed point in the usual regime $v^{2}\leq\frac{1}{2}$. Furthermore, if $mn>4$, the asymptotic velocity will satisfy $v^{2}>\frac{1}{2}$, regardless of whether there is a scaling regime or not, and we have provided evidence in Section 4.1.2 that $k(v)<0$ in this regime. All of this suggests the existence of a fixed point with $v^{2}>\frac{1}{2}$ in this case. To study the general case, it is useful to perform a full dynamical system analysis of the two evolution equations, as carried out in the next section (see Appendix C for more details). Before we continue, let us now make a brief comment on possible scaling solutions, i.e. stable (attractor) solutions with constant $v=\bar{v}$ and $L/t=\bar{\xi}$. As already noticed in Yamaguchi:2005gp , there is one crucial physical difference with respect to the constant tension case: the overall energy density in cosmic strings now scales as
$\rho=\frac{\mu}{L^{2}}=\frac{\mu_{0}t_{0}^{m}}{\xi^{2}t^{2+m}},$ (58)

### v2

@@ 3.3 Loop decays
In order to maintain the scaling regime, the network must lose energy at the same rate as the background. This is done by emission of loops. These loops then decay into axions, massive matter or gravitational waves, depending on the origin of the defect. We mentioned the discrepancies in the literature regarding these decay channels on the introduction, and here we will focus in local, Nambu-Goto strings. Simulations [96] suggest that around $90\%$ of the energy of the network is released into small ultra-relativistic loops that are red-shifted away. The other fraction $\mathcal{F}_{\alpha}\simeq 0.1$ of the energy is deposited into small loops of length that is a fraction $\alpha$ of the correlation length:
$\ell(t_{i})=\alpha L(t_{i})=\alpha\xi t_{i}\,,$ (3.15)
where in the second equality we have assumed the scaling regime of the VOS. Local cosmic string loops decay through gravitational radiation. [^14] Global strings can also decay by radiating massless Nambu-Goldstone bosons, with $\frac{dE_{\text{\tiny{GW}}}}{dt}\propto v^{2}$, where $v$ is the vev of the corresponding scalar field. In the frequency regime of interest, the power of emission is independent of their length, and reads
$\frac{dE_{\text{\tiny{GW}}}}{dt}=\Gamma G\mu^{2}\,,$ (3.16)
@@
$\ell(t)=\alpha t_{i}-\Gamma G\mu(t-t_{i})\,.$ (3.17)
Let us assume that the network loses energy at a rate [2]
$\frac{d\rho}{dt}=\tilde{c}\bar{v}\frac{\rho}{L}\,,$ (3.18)
@@
$\mathcal{F}_{\alpha}\frac{d\rho}{dt}=\mu\gamma\ell\frac{dn}{dt}\to\frac{dn}{dt}=\frac{\mathcal{F}_{\alpha}}{\gamma}\frac{\tilde{c}\bar{v}}{\alpha}\frac{1}{L^{4}}\,.$ (3.19)
where $\gamma$ is a Lorentz factor which takes into account that $\ell$ is the proper length of the loop, while the energy is computed in the rest frame of the mother string. The loop production rate during scaling therefore features a behaviour $dn/dt=A/t^{4}$ during scaling, where $A$ depends on the background [97, 96] and can be inferred from simulations.
@@ 4 The effect of a time-dependent tension
$\mu=m_{s}^{2}=\left(\frac{M_{P}}{4\pi}\right)^{2}\frac{\sqrt{g_{s}}}{\mathcal{V}},$ (4.1)
and the extra-dimensional volume $\mathcal{V}$ is measured in stringy units of $(2\pi\sqrt{\alpha^{\prime}})^{6}$. In phenomenologically interesting scenarios, such as volume kination or tracker epochs [66], the volume can vary with time, and thus induce a time-dependent string tension. In both cases the tension evolves as $\mu\sim t^{-m}$, with $m=1$ and $m=2/3$ for kination and the tracker respectively. The goal of this section will be to extend the original VOS model with a time-dependent string tension, and also discuss the evolution of loops151515 An early attempt to study the VOS with varying tension [29] imposed the string tension to depend on coordinate, rather than worldsheet time, in the Nambu-Goto action. This breaks worldsheet reparametrisations and leads to the wrong equations of motion (see also [34, 35]). It would be interesting to adapt the methods of [33] to the equations presented in this article.. We shall see that allowing for a varying tension entails a rich phenomenology, making scaling far from obvious. The goal of this article is to argue that a scaling solution is nevertheless achieved, and we will make some comments on the timescales for its stability to the percolation phenomenon recently discussed in [35]. Towards the end of the section we will also extend the analysis to the population of loops, whose distribution is crucial for the computation of gravitational wave spectra.
@@ 4.1 Equations of motion and analytical aspects
$S=-\int d^{2}\zeta\,\mu\left(X^{\nu}(\zeta)\right)\sqrt{-\gamma}.$ (4.2)
@@
$\frac{1}{\mu\left(X^{\nu}\right)\sqrt{-\gamma}}\partial_{a}\left(\mu\left(X^{\nu}\right)\sqrt{-\gamma}\gamma^{ab}g_{\rho\sigma}X^{\sigma}_{\,,b}\right)-\frac{1}{2}\partial_{\rho}g_{\lambda\sigma}\gamma^{ab}X^{\lambda}_{\,,a}X^{\sigma}_{\,,b}-\frac{\partial_{\rho}\mu\left(X^{\nu}\right)}{\mu\left(X^{\nu}\right)}=0,$ (4.3)
where the dependence on $X^{\nu}(\zeta)$ is left implicit. Eq. (4.3) was previously derived in [34, 35], and differs from the equation in [29] only by the last term, which arises precisely through the dependence of the tension on the embedding coordinates. From now onwards, we also assume that the tension only depends on time, that is $\mu\equiv\mu\left(X^{0}(\zeta)\right)$. As in the case of constant tension, we choose a gauge where $\dot{X}\cdot X^{\prime}=0$ (and $\gamma_{01}=\gamma_{10}=0$), and identify conformal and worldsheet time, $X^{0}(\zeta)=\tau$. Notice, however, that this is only possible after the equations of motion have been written down out, as doing so in the action would have lead to a wrong result for (4.3). Defining the energy density of a string as in (3.4), the equation of motion (4.3) for $\rho=0$ obtains an additional term with respect to the constant tension case
$\dot{\epsilon}+\left(2\frac{\dot{a}}{a}+\frac{\dot{\mu}}{\mu}\right)\varepsilon\dot{x}^{2}=0.$ (4.4)
@@
$\ddot{x}_{i}+\left(2\frac{\dot{a}}{a}+\frac{\dot{\mu}}{\mu}\right)(1-\dot{x}^{2})\dot{x}_{i}-\frac{1}{\varepsilon}\left(\frac{x_{i}^{{}^{\prime}}}{\varepsilon}\right)^{{}^{\prime}}=0,$ (4.5)
@@
$2\mathcal{H}^{\prime}=2\mathcal{H}+\frac{\dot{\mu}}{\mu}.$ (4.6)
@@
The interplay between varying tension and Hubble expansion leads to a rich phenomenology. For instance, as explained in [34], for the special case $\mu\sim a(t)^{-2}$ the action becomes scale-invariant, and the equations of motion reproduce those of a constant tension string in flat space-time. Even more striking is the case when the RHS of (4.6) is negative, since the loops grow faster than the scale factor [35]; the strings behave as if they were placed in a contracting universe. It is this case that we intend to study in this article [^17] Note that the universe is still undergoing expansion, so the dynamics of the network will be different to that of cosmic strings in an actual contracting universe, studied in [98, 99]..
@@ 4.1.1 The circular loop
To be more concrete about our claims, let us find a solution for the microscopic equations of motion in an arbitrary power-law cosmology $a\sim t^{2/n}$ with $\mu\sim t^{-m}$. Note that this includes the constant tension case $m=0$, but we are not aware of this solution having been reported elsewhere in the literature (although its qualitative features are understood). It is well known that, when $\textbf{x}^{\prime}\cdot\dot{\textbf{x}}=0$, the equations of motion in Eqs. (4.4) and (4.5) are not independent. To solve the condition $\textbf{x}^{\prime}\cdot\dot{\textbf{x}}=0$ we focus on the circular loop:
$\textbf{x}=R(\tau)L(\cos(\zeta^{1}/L),\sin(\zeta^{1}/L),0)\,,$ (4.7)
@@
$R_{max}(\tau)\sim\left(\frac{\tau}{\tau_{0}}\right)^{-q/2}\,,$ (4.9)
@@ 4.1.2 The momentum parameter: analytical modelling
The momentum parameter $k(v)$ plays a crucial role in the averaged equations of motion, and in this section we would like to gain some understanding on its behaviour from a microscopic point of view. As will be clear in the next section, a crucial aspect of achieving a scaling regime (when $m\,n>4$) is that the momentum parameter switches sign when crossing the threshold $v^{2}=\frac{1}{2}$. Let us provide some evidence in support of this claim. To do so, we study the evolution of small scale structure with varying tension. Let us follow [4] and study linear perturbations with respect to the static, straight string:
$\textbf{x}=\textbf{c}\zeta^{1}\,,$ (4.10)
@@
$\delta\ddot{x}_{i}+\left(2\frac{\dot{a}}{a}+\frac{\dot{\mu}}{\mu}\right)\delta\dot{x}_{i}-\delta x_{i}^{\prime\prime}=0\,.$ (4.11)
@@
$\delta x=\textbf{A}_{k}\left(\frac{\tau_{0}}{\tau}\right)^{\nu}J_{\nu}(k\tau)e^{ik\zeta^{1}}+\textbf{B}_{k}\left(\frac{\tau_{0}}{\tau}\right)^{\nu}Y_{\nu}(k\tau)e^{ik\zeta^{1}}\,,$ (4.12)
@@
$\lim_{k\tau\to\infty}J_{\nu}(k\tau)=\sqrt{\frac{2}{\pi k\tau}}\cos\left(z-\frac{\nu\pi}{2}-\frac{\pi}{4}\right)\,,$ (4.13)
@@
$\langle(1-\dot{\textbf{x}}^{2})\dot{\textbf{x}}\cdot\textbf{u}\rangle=\langle\frac{R}{a}\frac{1}{\epsilon^{2}}\textbf{x}^{\prime\prime}\cdot\dot{\textbf{x}}\rangle\,.$ (4.14)
@@
$\langle\frac{R}{a}\frac{1}{\epsilon^{2}}\textbf{x}^{\prime\prime}\cdot\dot{\textbf{x}}\rangle=\frac{1}{16a}\textbf{A}_{k}^{2}k^{2}\left(\left(\frac{\tau_{0}}{\tau}\right)^{2\nu}J_{\nu}(k\tau)^{2}-\frac{1}{2^{\nu}\Gamma(1+\nu)^{2}}\right)\,,$ (4.15)
where $\Gamma(x)$ is Euler’s gamma function and we have averaged, for simplicity, from an initial time $\tau=0$. The crucial point here is that for subhorizon perturbations, using Eq. (4.13), the first term dominates for $02\nu+1<$ 0, which is equivalent to $mn<4$. Otherwise the second term dominates and the contribution to the momentum parameter is negative - we take this as suggestive evidence of a flip of sign in the momentum parameter for the VOS model when $mn>4$ and therefore for the existence of scaling solutions in this regime, as explained below. We believe this is so even if our solutions do not fully capture the complicated microscopic configuration of the long strings in a scaling solution. In particular, the microscopic equations cannot take into account the loss of energy into loops by interactions of the strings, which is a key ingredient for scaling to occur.
@@ 4.2 Long strings
In this section, we examine how the modified equations of motion (4.4)-(4.5) will impact the evolution and properties of the network, focusing on the population of the long strings.
@@ 4.2.1 Network evolution
From (4.4), the energy density $\rho\sim\frac{E}{a^{3}}$ of strings then obeys the equation
$\frac{{\rm d}\rho}{{\rm d}t}=-2H(1+v^{2})\rho+\frac{1}{\mu}\frac{{\rm d}\mu}{{\rm d}t}(1-v^{2})\rho$ (4.16)
where we have switched to the FRW time coordinate $t$. Notice that, in the non-relativistic limit $(v\rightarrow 0)$, the energy density redshifts as $\rho\sim\mu(t)/a(t)^{2}$, while in the ultra-relativistic limit $(v\rightarrow 1)$ it behaves as radiation, $\rho\sim 1/a(t)^{4}$. From the definition of the characteristic length $\rho=\frac{\mu}{L^{2}}$, one can turn (4.16) into an equation for $L$. For long strings, one also has to take into account energy loss into loops, which can be modelled as
$\frac{{\rm d}\rho}{{\rm d}t}=-\tilde{c}v\frac{\rho}{L}.$ (4.17)
Putting everything together, (4.16)-(4.17) can be combined to give
$2\frac{{\rm d}L}{{\rm d}t}=2HL(1+v^{2})+\tilde{c}v+\frac{L}{\mu}\frac{{\rm d}\mu}{{\rm d}t}v^{2}.$ (4.18)
Following the same steps, one can use (4.5) to derive a new equation for $v$, namely
$\frac{{\rm d}v}{{\rm d}t}=(1-v^{2})\left[\frac{k(v)}{L}-\left(2H+\frac{1}{\mu}\frac{{\rm d}\mu}{{\rm d}t}\right)v\right].$ (4.19)
The difference between these and those of [33] were explained in footnote 15. Here, one can already see a crucial, qualitative difference with respect to the constant tension case: if $02H+\frac{1}{\mu}\frac{{\rm d}\mu}{{\rm d}t}<$ 0, the RHS of (4.19) cannot be negative for any $v$ satisfying $k(v)>0$, forbidding the existence of a fixed point in the usual regime $v^{2}\leq\frac{1}{2}$. Furthermore, if $mn>4$, the asymptotic velocity will satisfy $v^{2}>\frac{1}{2}$, regardless of whether there is a scaling regime or not, and we have provided evidence in Section 4.1.2 that $k(v)<0$ in this regime. All of this suggests the existence of a fixed point with $v^{2}>\frac{1}{2}$ in this case. To study the general case, it is useful to perform a full dynamical system analysis of the two evolution equations, as carried out in the next section (see Appendix C for more details). Before we continue, let us now make a brief comment on possible scaling solutions, i.e. stable (attractor) solutions with constant $v=\bar{v}$ and $L/t=\bar{\xi}$. As already noticed in [29], there is one crucial physical difference with respect to the constant tension case: the overall energy density in cosmic strings now scales as
$\rho=\frac{\mu}{L^{2}}=\frac{\mu_{0}t_{0}^{m}}{\xi^{2}t^{2+m}},$ (4.20)